OpenStudy (anonymous):

Find the inverse laplace transform of 1/((s^2+1)(s^2+2s+3)).

zepdrix (zepdrix):

Hmm I'm not really sure how to approach this one. Could we do this maybe?\[\large s^2+2s+3 \qquad=\qquad (s^2+2s+1)+2 \qquad=\qquad (s+1)^2+2\] Then I'm not sure if I'm setting up these partial fractions up correctly, or if they'll even work. Just an idea though:\[\large \frac{1}{(s^2+1)\left[(s+1)^2+2\right]}=\frac{As+B}{s^2+1}+\frac{C(s+1)+D}{(s+1)^2+2}\]

OpenStudy (anonymous):

Let me try.

OpenStudy (anonymous):

How do you solve for A, B, C, D?

zepdrix (zepdrix):

Expand eeeeeeeverything out. :[ Equate like terms.

OpenStudy (loser66):

back to your previous post to get it, girl

zepdrix (zepdrix):

Oh did you mean to type:\[\Large s^2+3s+2\]Idealist? Because that's quite a bit different than what you have above,\[\Large s^2+2s+3\]:(

OpenStudy (anonymous):

@zepdrix \[\frac{1}{(s^2+1)\left[(s+1)^2+2\right]}=\frac{As+B}{s^2+1}+\frac{C\color{red}s+D}{(s+1)^2+2}\] Using \(C(s+1)\) gives you \(Cs+C\), which mucks up the next few steps. You'll find that \(D=0\), whereas using \(Cs\) gives you \(D=\dfrac{1}{4}\).